# TRACE Back from the Future: A Probabilistic Reasoning Approach to Controllable Language Generation

### v1

@@ 5.1 Experimental Setup
Evaluation Metrics.
Metrics vary by task. For the primary detoxification task (Section 5.2), we follow the setup in Liu et al. (2021) using the RealToxicityPrompts dataset (Gehman et al., 2020). We evaluate Toxicity (Perspective API avg. max. toxicity & prob. of any toxic generation ($>0.5$) over 25 samples; ↓ lower is better), Perplexity (as an automatic measure of fluency, calculated using GPT2-XL; ↓ lower is better), and Diversity (Distinct 2-grams, 3-grams; ↑ higher is better). We also include supplementary AI evaluations using GPT4o-mini. For role-playing (Section 5.3), we measure role quality via classifier probability. For topic control (Section 5.5), we measure political relevance using the zero-shot classifier from Laurer et al. (2023). Full details on metrics are provided in Appendix E.
Baselines.
TRACE is compared against representative baselines including fine-tuning (DAPT), RL (PPO, Quark, DPO), and various decoding methods (PPLM, GeDi, FUDGE, DExperts, MuCoLa). Results are sourced primarily from prior work (Liu et al., 2021; Lu et al., 2022; Kumar et al., 2022) using the same setup as in Liu et al. (2021) or run by us for direct comparison (DPO, prompting). Full details on baseline implementations and sourcing are provided in Appendix D.
Implementation Details.
The HMMs used by TRACE were distilled [^2] The distillation process involves compiling a HMM to an equivalent probabilistic circuit (Choi et al., 2020) on the GPU, which is trained on LM samples using the expectation-maximization (EM) algorithm (Dempster et al., 1977). from base LMs (GPT2-Large, Gemma-2B) following Zhang et al. (2023); Liu et al. (2024b); specific hyperparameters and training details are in Appendix B. Once distilled, the HMM model is fixed and is reused without further training for each attribute. Attribute classifiers (e.g., for non-toxicity, non-politicalness) were trained according to the methodology in Section 4.3; specific datasets, oracles, fitting procedures (log-MSE), and transformation parameters ($b,c$) are detailed in Appendix C.
@@ 5.2 Detoxification
Figure 2: Effect of Probability Transformations on Distribution of Non-toxicity. (a) Training-time transformation creates a bimodal target distribution, concentrating probability near 0 (toxic region). This assists log-MSE loss (Sec 4.3) in learning stronger penalties for undesired tokens, aiding classifier learning, though the fitted distribution may remain unimodal. (b) Decoding-time transformation reshapes the computed EAP distribution towards a bimodal form during generation, enhancing attribute control via sharper separation of compliant/non-compliant futures.
In this experiment, we benchmark TRACE for the task of generating nontoxic continuations for prompts from the RealToxicityPrompts dataset (Gehman et al., 2020). Our method, TRACE, generally outperforms baselines in detoxification while maintaining competitive fluency and diversity, as shown in Table 1. Specifically, TRACE surpasses both finetuned/RL models and sampling- or discriminator-based decoding methods in achieving low toxicity with high language quality.
We note that while some baseline methods, particularly those based on RL fine-tuning, exhibit lower absolute perplexity than the baseline GPT-2 (indicated by struck-through scores in Table 1), they also typically suffer from extremely low n-gram diversity. Further supporting this observation, the DPO baseline shows significantly lower generation entropy compared to the base model and TRACE (Table 2). As observed by Holtzman et al. (2020), excessively low perplexity, especially when coupled with low diversity and entropy, often signals repetitive or unnatural text rather than improved language quality. Therefore, we consider the original GPT-2 model a strong perplexity baseline. Since detoxification inherently constrains the generation space, a slight increase in perplexity relative to the unconstrained baseline is expected for methods that effectively perform detoxification while maintaining diversity. Among the evaluated methods that maintain reasonable diversity (comparable to GPT-2), TRACE achieves the most favorable perplexity (closest to the baseline) while preserving high n-gram diversity scores.
Method Entropy (↑)
GPT2-large 52.06
DPO 39.52
TRACE 52.54
Table 3: GPT4 LM-as-a-judge Evaluations.
Method Nontoxicity (↑) Fluency (↑) Diversity (↑)
Gemma2B 4.39 3.76 2.93
DPO 4.65 3.94 2.86
TRACE 4.69 3.72 2.94
Factorized Classifier is Effective Despite Attributes Being Non-Factorizable. We compared the classification performance of our factorized classifiers against neural classifiers (details in Appendix A). A clear gap in performance, measured by cross-entropy against oracle scores, confirms that attributes like toxicity are not fully factorizable. Despite this inherent limitation of the factorized classifier, TRACE demonstrates superior performance in controllable generation tasks compared to approaches that utilize more complex neural classifiers but rely on approximations of the Expected Attribute Probability (EAP), such as GeDi (Krause et al., 2021) and FUDGE (Yang & Klein, 2021). This result suggests, perhaps surprisingly, that exact EAP computation (using an HMM), even when paired with a simpler, appropriately trained factorized classifier, can be as effective, or even more so, than using more complex classifiers within less exact EAP estimation frameworks.
Scaling to Larger LMs. To evaluate the scalability and effectiveness of TRACE on larger models, we distilled an HMM from Gemma-2B (Team et al., 2024) and applied our method. For comparison, we also trained DPO, a RL baseline, on Gemma-2B, which displays strong detoxification at the cost of reduced diversity. As shown by both automatic metrics (Table 1) and GPT-4 evaluations (Table 3), TRACE achieves superior detoxification compared to DPO on Gemma-2B, while better maintaining both fluency and diversity close to the base model. These results are comparable to the performance gains observed with GPT2-large, indicating that TRACE effectively enhances controllability for larger, modern language models.
@@
Training- and Decoding-Time Transformations Enhance Attribute Satisfaction. The logit-space probability transformation (defined in Section 4.3) offers complementary benefits when applied at different stages of TRACE, enhancing overall attribute control.
Applying the transformation during classifier training reshapes the target probability distribution used for log-MSE fitting (Section 4.3). As illustrated in Figure 2 (top, Tranformed Distribution), this can create a more bimodal target. Effectively, this helps concentrate the target distribution towards the low-probability region corresponding to rare attributes (e.g., toxicity, when fitting a non-toxicity model), assisting the log-MSE objective in learning appropriately strong penalties (low weights $l(x_{i})$) for the associated tokens. While this aids the learning process, we empirically observe that the fitted classifier distribution often remains unimodal (Figure 2 (top, Fitted Distribution)).
The desired bimodal probability distribution for effective control is instead achieved by applying the transformation during decoding. This modifies the computed EAP $p_{\textit{hmm}}(s\mid x_{t},x_{<t})$ at each step before it influences generation (Eq. 9). As seen in Figure 2 (bottom), this directly reshapes the EAP distribution, turning a potentially diffuse or unimodal EAP into a sharper, bimodal one. This enhances the separation between attribute-compliant and non-compliant continuations, leading to stricter enforcement during generation.
Thus, the two applications of transformation are complementary: training-time transformation facilitates better learning of attribute weights via log-MSE, while decoding-time transformation provides more direct control over the generation by shaping the EAP distribution for stricter compliance.
Decoding Transformation Enables Fluency-Attribute Trade-off Another key advantage of applying the transformation during decoding is that its scaling parameter $b$ acts as an intuitive control knob. By adjusting this ’decoding multiplier’, users can directly modulate the strength of the attribute control during generation. As illustrated for detoxification in Figure 3, increasing $b$ enhances attribute compliance (lower toxicity) but may decrease fluency (higher perplexity). This provides a straightforward mechanism to customize the trade-off between fluency and attribute satisfaction based on specific application requirements, without the need for any retraining.
Figure 3: Decoding Time Transformation Enables Trading Off Fluency for Detoxification. As the decoding multiplier increases from 1 to 1.8, fluency decreases while detoxification improves, demonstrating the ability to balance these two aspects by adjusting a single hyperparameter.
Impact of HMM Quality Our results indicate that the quality of the distilled HMM significantly impacts TRACE’s performance. To investigate this, we trained two HMMs for Gemma-2B using identical hyperparameters but different amounts of training data: one with 1 million samples, denoted TRACE (HMM), and another with 10 million samples, used in our primary TRACE results (details in Appendix B). As shown in Table 1, the TRACE (HMM) configuration underperforms the standard TRACE setup on detoxification metrics. This comparison highlights that increased data volume for HMM training directly improves downstream controllable generation effectiveness. Consequently, further advancements in HMM distillation techniques or using larger training datasets may yield even better performance from TRACE.
@@ 5.3 Role-playing 76 Characters with Limited Data
Figure 4: Personalized LM: TRACE outperforms prompting engineering in role quality.
Figure 5: Training Time vs. Quality for Role Imitation. The scatter plot illustrates training time and quality scores for 76 different characters using TRACE. Training duration clusters around three seconds per new attribute, achieving high-quality results even in low-data scenarios.
Lightweight and Low-Resource Adaptation. A primary advantage of TRACE is its ability to integrate new attributes by training only a lightweight log-linear classifier. This facilitates rapid adaptation and supports attributes with limited data, addressing the need for low-resource personalized LMs. Leveraging this flexibility, we personalized both GPT2-large and Gemma-2B for 76 distinct characters sourced from the RoleBench dataset (Wang et al., 2024). For each character, the TRACE classifier was trained using the RoleBench training split (approx. 300 question-answer pairs per role).
Table 4 qualitatively compares example responses generated by Gemma-2B for three characters. This comparison includes the base model’s direct response (Direct Prompt), the base model guided by TRACE, and a few-shot prompting baseline. This prompting baseline uses a standard role instruction plus 10 QA examples specific to the character, sampled from the RoleBench training set (details in Appendix D). Qualitatively, the TRACE-guided answers exhibit distinct content and tone reflective of the characters’ personalities compared to the baselines.
Quantitatively, evaluating on GPT2-large, TRACE achieves superior role quality compared to prompting across the vast majority of the 76 characters (Figure 4). This evaluation compares TRACE against the Role-Specific Instruction prompting baseline (details in Appendix D). We use this simpler prompting setup here to specifically assess performance based on instruction following alone, without few-shot examples. This provides a direct comparison where both TRACE and the baseline operate with minimal input overhead, in contrast to few-shot prompting which incurs higher decoding costs (see Section 5.4, Table 6).
Figure 5 (using GPT2-large) further highlights TRACE’s efficiency for this low-resource task: adapting to each new character requires only about three seconds of classifier training time on average, while achieving the high role quality scores shown. This underscores TRACE’s capability for rapid personalization even with minimal data per attribute.
@@ 5.4 Training and Inference Time Analysis
Table 5: Training Time for Each New Attribute
Method Training Time
Mix and Match 2 hours
DExperts 3 min–16 hours
DAPT 16 hours
GeDi 5 hours
TRACE 10 seconds
Table 6: Inference Time Relative to Baseline
Method Inference Ratio
@@
TRACE 1.1
Table 7: Composition results (Political+Nontoxic) on RealToxicityPrompts with GPT2-Large. Metrics: Tox (avg. max./prob.$>$0.5; ↓), Mean Pol ↑, PPL ↓, Diversity (Dist-2/3; ↑). """"+ Dec. TF"""" applies decoding TF (Sec 4.3) only to the Political EAP.
Models Max Tox (↓) Any Tox $>$ 0.5 (↓) Mean Pol (↑) PPL (↓) Dist-2 (↑) Dist-3 (↑)
@@
TRACE (Detox + Pol + Dec. TF) 0.187 0.024 0.516 32.78 0.87 0.86
TRACE is designed for rapid adaptation to new attributes while maintaining efficient inference compared to alternative approaches. The framework requires a one-time HMM distillation from the base language model, independent of specific control attributes.
Once the HMM is trained, adapting TRACE to a new attribute requires only fitting a lightweight log-linear classifier. As demonstrated in the role-playing experiments (Figure 5), this process is extremely fast for low-resource attributes, taking about three seconds per character on average. Even for attributes trained on larger datasets (tens of thousands of samples), fitting takes only up to ten seconds (Table 5). This contrasts sharply with methods requiring extensive retraining or auxiliary model training; for context, previously reported training times for baselines like DAPT, DExperts, GeDi, and Mix and Match range from minutes to over 16 hours on various GPUs (Table 5). This rapid adaptation makes TRACE highly suitable for dynamic or low-resource scenarios.
During generation, TRACE maintains efficiency. The HMM’s expected attribute probabilities (EAP) involve a precomputed backward pass (Section 4.2) and an efficient forward update per token. We implement the HMM computations on the GPU alongside the base language model. For single-sequence (non-batched) generation, this introduces only a minor overhead, resulting in approximately $\times1$ the inference time compared to the baseline language model (Table 6). The overall generation complexity remains dominated by the base LM, scaling linearly with sequence length $l$ (i.e., $\mathcal{O}(l)$ assuming constant per-token LM cost). This efficiency stands in contrast to methods involving lengthy prompts for in-context learning, external discriminators (GeDi, DExperts), or iterative sequence updates (Mix and Match, MuCoLa, PPLM), which incur substantially higher decoding costs (Table 6).
@@ 5.5 Composition: Political and Nontoxic Texts
A key benefit of TRACE is its ability to compose multiple attributes without retraining, which is especially advantageous for rare or novel attribute combinations. Given two attributes $s_{1},s_{2}$ with corresponding classifiers, TRACE enables conditioning on their conjunction by multiplying their probabilities during decoding: $p(s_{1}\text{ AND }s_{2}|x_{1:n})=p(s_{1}|x_{1:n})p(s_{2}|x_{1:n})$. This composition relies on the modeling assumption that the attributes are independent given the text $x_{1:n}$, and naturally aligns with TRACE’s factorized classifiers, as the combined probability corresponds to a new factorized classifier with weights $w^{\prime}=w^{1}\cdot w^{2}$.
To illustrate compositionality, we generated texts that are simultaneously political and nontoxic. While perhaps intuitively anticorrelated, these attributes were not found to be significantly statistically anticorrelated within the RealToxicityPrompts (RTP) dataset. Nonetheless, their combination presents a challenging control task due to data scarcity for such composite attributes, often making methods requiring specific joint training infeasible. TRACE, requiring only individually trained classifiers, remains effective.
Since combining attributes is challenging, most prior work does not explore this scenario. We conduct ablation experiments (Table 7) using GPT2-large on 7500 RTP prompts to demonstrate TRACE’s effectiveness. The results demonstrate the effectiveness of composition: controlling for only a single attribute (e.g., ‘TRACE (NonToxic only)‘ or ‘TRACE (Political only)‘) primarily affects that specific attribute, leaving the other largely unchanged relative to the baseline. In contrast, the compositional approach (‘TRACE (NonToxic + Political)‘) successfully reduces toxicity to levels comparable to the ‘NonToxic only‘ setting, while simultaneously increasing political relevance to levels comparable to the ‘Political only‘ setting.
@@ 6 Conclusion
We introduced TRACE, a lightweight and adaptable framework for controllable text generation leveraging tractable probabilistic inference. The core strengths of TRACE lie in its efficiency, adaptability, and strong performance achieved without modifying or finetuning the base LM. By distilling a Hidden Markov Model (HMM) from a base LM and combining it with simple, efficiently trained attribute classifiers, TRACE enables tractable computation of Expected Attribute Probability (EAP) to guide generation towards desired attributes. Empirically, TRACE achieves state-of-the-art detoxification performance with minimal decoding overhead ($\sim$10%). Furthermore, it demonstrates remarkable efficiency and effectiveness in low-resource scenarios, adapting to personalize generation for 76 distinct characters in seconds, and successfully composes multiple attributes, such as generating text that is simultaneously political and non-toxic.
Despite the strong empirical performance of TRACE, there is potentially scope to further boost performance by improving the expressivity of both the HMM language model and the attribute classifier. Our results show that HMMs trained on more data yield better results (Section 3), suggesting future improvements in HMM distillation techniques could further enhance performance. Additionally, while we show that using factorized classifiers is effective even for attributes that are not perfectly factorizable, extending TRACE to incorporate more expressive classifiers (e.g., based on tractable probabilistic circuits (Khosravi et al., 2019; Choi et al., 2020)) that remain compatible with efficient EAP computation presents a promising avenue for future work, potentially enabling stronger control over complex semantic attributes like long-range coherence.

### v2

@@ 5.1 Experimental Setup
Evaluation Metrics. Metrics vary by task. For the primary detoxification task (Section 5.2), we follow the setup in Liu et al. (2021) using the RealToxicityPrompts dataset (Gehman et al., 2020). We evaluate Toxicity (Perspective API avg. max. toxicity & prob. of any toxic generation (>0.5>0.5) over 25 samples; ↓ lower is better), Perplexity (as an automatic measure of fluency, calculated using GPT2-XL; ↓ lower is better), and Diversity (Distinct 2-grams, 3-grams; ↑ higher is better). We also include supplementary AI evaluations using GPT4o-mini. For role-playing (Section 5.3), we measure role quality via classifier probability. For topic control (Section 5.5), we measure political relevance using the zero-shot classifier from Laurer et al. (2023). Full details on metrics are provided in Appendix F.
Baselines. TRACE is compared against representative baselines including fine-tuning (DAPT), RL (PPO, Quark, DPO), and various decoding methods (PPLM, GeDi, FUDGE, DExperts, MuCoLa). Results are sourced primarily from prior work (Liu et al., 2021; Lu et al., 2022; Kumar et al., 2022) or run by us (DPO) using the same setup as in Liu et al. (2021). Full details are in Appendix E.
Implementation Details. The HMMs used by TRACE were distilled [^3] The distillation process involves compiling a HMM to an equivalent probabilistic circuit (Choi et al., 2020) on the GPU, which is trained on LM samples using a mini-batch variant of the expectation-maximization (EM) algorithm (Dempster et al., 1977). from base LMs (GPT2-Large, Gemma-2B) following Zhang et al. (2023); Liu et al. (2023); details on the different HMM configurations, including hidden state sizes and training parameters, are in Appendix C. Once distilled, the HMM model is fixed and is reused without further training for each attribute. Attribute classifiers (e.g. nontoxicity) were trained according to the method in Section 4.3; datasets, oracles, fitting procedures, and transformation parameters ($b,c$) are detailed in Appendix D.
@@ 5.2 Detoxification
We evaluate TRACE on detoxification using RealToxicityPrompts (Gehman et al., 2020), comparing performance against fine-tuning (DAPT), RL (PPO, Quark, DPO), and decoding methods (GeDi, FUDGE, DExperts, MuCoLa).
TRACE Outperforms Existing Methods Table 1 shows TRACE significantly reduces toxicity compared to baselines, and maintains high diversity, with minor reductions from GPT-2’s baseline scores. RL methods (e.g., PPO, Quark, DPO), despite lowering toxicity, sharply reduce diversity and perplexity, indicative of repetitive or unnatural text generation due to mode collapse. Appendix Tables 7 and 8 further support these observations: TRACE achieves higher conditional entropy than DPO, and receives comparable or better ratings from GPT-4 on nontoxicity and diversity.
Factorized Classifier is Effective Despite Non-Factorizable Attributes.
Attributes like toxicity are not fully factorizable, as evidenced by the performance gap between neural and our factorized classifiers (Appendix D). Nonetheless, TRACE outperforms methods like GeDi and FUDGE that use more expressive neural classifiers but rely on approximate EAP estimation. This highlights a key insight: exact EAP computation over an HMM, even with a simple factorized classifier, can rival or surpass more complex classifiers when the inference framework is less precise.
Scaling TRACE to Larger Language Models. TRACE scales effectively to larger language models such as Gemma-2B. As shown in Table 1 and supported by GPT-4 evaluations (Appendix Table 8), TRACE consistently outperforms DPO—a reinforcement learning-based method—in detoxification performance, while maintaining strong fluency and diversity. These results demonstrate TRACE’s broad applicability and robustness across model scales.
Table 2: Ablation study of TRACE variants on GPT-2 Large Rows: no transformation, training-time transformation only, or both training- and decoding-time transformation.
TRACE Variant Toxicity (↓) Diversity (↑) PPL (↓)
avg. max. prob. dist-2 dist-3
No Transformation 0.353 0.196 0.87 0.86 25.44
Training TF 0.187 0.026 0.87 0.85 27.51
Train + Dec TF 0.163 0.016 0.85 0.85 29.83
Figure 2: Complementary Effects of Probability Transformation. Top: At training time, the transformation concentrates oracle probabilities for undesirable attributes towards 0 to improve classifier learning. Bottom: At decoding time, it reshapes the unimodal EAP into a bimodal one to enforce stricter generation control.
Figure 3: Tuning the Fluency-Detoxification Trade-off. The decoding-time transformation multiplier acts as a control knob; as it increases, detoxification improves while fluency decreases.
Complementary Roles of Training and Decoding Transformations
The probability transformation offers complementary benefits when applied at different stages of TRACE, as shown by the ablation study in Table 2. Applying the transformation at training time alone (“Training TF”) provides an initial reduction in toxicity over the baseline TRACE model with no transformation. The most significant gain comes from applying the transformation at both stages (“Train + Dec TF”); this further reduces average maximum toxicity from 0.187 to 0.163, albeit with a corresponding increase in perplexity. Figure 2 visually explains the complementary mechanisms driving these gains.
Top: At training time, the transformation creates a stricter learning target by reshaping the skewed Original Dist.” into a bimodal ”Transformed Dist.”, critically pushing scores for toxic content towards 0. The log-MSE objective’s focus on this low-probability region helps the classifier learn stronger negative weights for toxic tokens, resulting in a ”Fitted Dist.” that successfully captures this new peak near 0.
Bottom: At decoding time, the same transformation is applied directly to the computed Expected Attribute Probability (EAP). It converts a diffuse, unimodal EAP distribution (“Before Trans.”) into a sharp, bimodal one (“After Trans.”), creating a clearer separation between attribute-compliant and non-compliant continuations.
@@
This synergy—where one transformation aids learning and the other enforces control—accounts for their strong combined performance. As a final benefit, the decoding transformation’s scale parameter ($b$) offers an intuitive knob for tuning this control, allowing users to smoothly trade toxicity for fluency post-hoc (Figure 3).
Impact of HMM Quality Finally, the effectiveness of control is sensitive to the quality of the HMM itself. As shown in Appendix Fig. 5, as the distilled HMM’s log-likelihood improves with more training, we observe a corresponding decrease in generation toxicity. This highlights that better HMM distillation directly enhances TRACE’s performance.
Figure 4: TRACE outperforms prompt engineering in role quality.
@@ 5.3 Role-playing 76 Characters with Limited Data
Lightweight and Low-Resource Adaptation. A key advantage of TRACE is its rapid, low-resource adaptation; new attributes are integrated by training only a lightweight log-linear classifier. Leveraging this, we personalized both GPT2-large and Gemma-2B for 76 distinct characters from the RoleBench dataset (Wang et al., 2024). Each character’s classifier was trained on its corresponding RoleBench training split of $\sim$ 300 question-answer pairs.
Table 3 qualitatively compares Gemma-2B responses for three characters across three settings: the base model, a few-shot prompting baseline, and TRACE. The prompting baseline uses a role instruction plus 10 question-answer examples from the RoleBench training set (details in Appendix E). Qualitatively, TRACE-guided answers show more distinct, character-reflective content and tone than the baselines.
Quantitatively, on GPT2-large, TRACE achieves superior role quality over a standard prompting baseline for most of the 76 characters (Fig. 4) . This baseline uses a role-specific but exemplar-free instruction (details in Appendix E), ensuring a direct comparison of instruction-following with minimal input overhead, unlike few-shot prompting which has higher decoding costs (see Section 5.4, Table 4).
@@ 5.4 Training and Inference Time Analysis
Table 4: Training and inference time comparison. TRACE achieves minimal training overhead and near-baseline inference cost.
Method Train Time
Mix and Match 2 hr
DExperts 3 min–16 hr
DAPT 16 hr
GeDi 5 hr
TRACE 10 s
Method Inf. Ratio
@@
TRACE 1.2
Table 5: Composition results (Political+Nontoxic) on RealToxicityPrompts with GPT2-Large. Metrics: Tox (avg. max./prob.>>0.5; ↓), Mean Pol ↑, PPL ↓, Diversity (Dist-2/3; ↑). "+ Dec. TF" applies decoding TF (Sec 4.3) only to the Political EAP.
Models Max Tox (↓) Any Tox >> 0.5 (↓) Mean Pol (↑) PPL (↓) Dist-2 (↑) Dist-3 (↑)
@@
TRACE is designed for rapid adaptation and efficient inference, requiring only a one-time HMM distillation from a base LM, independent of any specific control attribute.
Training Time. Once the HMM is trained, adapting to a new attribute only requires fitting a lightweight classifier. This process takes up to 10 seconds, in sharp contrast to baselines like DAPT, DExperts, GeDi, and Mix and Match, which requires anywhere from minutes to 16 hours of training on various GPUs. This rapid adaptation makes TRACE highly suitable for dynamic or low-resource scenarios.
Inference Time. TRACE maintains efficiency during generation. While alternative approaches that rely on in-context learning, discriminators, or iterative updates incur substantial decoding costs—with inference ratios ranging from 7.5x to 40x the baseline—TRACE introduces only minor overhead. This efficiency stems from its design; the HMM’s Expected Attribute Probability (EAP) calculation uses a precomputed backward pass and an efficient forward update per token. Consequently, the total inference time is only about 1.2x that of the baseline language model, as the process remains dominated by the base model’s computation.
@@ 5.5 Composition: Political and Nontoxic Texts
A key benefit of TRACE is its ability to compose multiple attributes without retraining. Conditioning on the conjunction of two attributes ($s_{1},s_{2}$) is achieved by multiplying their probabilities during decoding, based on the modeling assumption of attribute independence given the text ($x$): $p(s_{1}\text{ and }s_{2}|x)=p(s_{1}|x)p(s_{2}|x)$. For TRACE’s factorized classifiers, this simplifies to creating a new composite classifier by multiplying the token weights ($w$) of the individual classifiers: $w^{\prime}=w^{1}\cdot w^{2}$. To demonstrate, we task the model with generating text that is simultaneously political and nontoxic. This combination is rare, making joint training challenging for other methods, but TRACE remains effective as it only needs the individually trained classifiers.
The results in Table 5 demonstrate effective composition. Controlling for a single attribute primarily affects only that dimension: TRACE (Detox only) reduces Max Toxicity from 0.386 to 0.186 while leaving the Mean Political score nearly unchanged (0.169 vs 0.168). Conversely, TRACE (Pol only) increases the Mean Political score from 0.169 to 0.333 with almost no change in toxicity. The compositional TRACE (Detox + Pol) approach successfully achieves both goals simultaneously, reaching a toxicity level (0.190) and a political score (0.344) comparable to their respective single-attribute control settings.
@@ 6 Conclusion
We introduced TRACE, a lightweight framework for controllable text generation that uses tractable probabilistic inference. Its core strengths are efficiency, adaptability, and strong performance, all achieved without modifying or finetuning the base LM. TRACE distills a Hidden Markov Model (HMM) from a base LM and combines it with simple, efficiently trained classifiers to tractably compute the Expected Attribute Probability (EAP), which guides generation towards desired attributes.
Empirically, TRACE achieves state-of-the-art detoxification with only a $\sim$20% decoding overhead. It also excels in low-resource scenarios, adapting to personalize generation for 76 distinct characters in seconds, and successfully composes multiple attributes, such as generating text that is simultaneously political and non-toxic.
Despite TRACE’s already strong empirical performance, its capabilities can be further boosted by improving the expressivity of both the HMM and the attribute classifier. Our results show that higher-quality HMMs yield better results, suggesting that improved distillation techniques are a promising direction. Additionally, while our simple factorized classifiers are effective even for non-factorizable attributes, extending TRACE to incorporate more expressive models—such as tractable probabilistic circuits (Khosravi et al., 2019; Choi et al., 2020)—that remain compatible with efficient EAP computation could enable stronger control over complex attributes like long-range coherence.